OpenStudy (anonymous):

Differentiate (2*n*b^(r*x)+n)^k. In this question when it asks to differentiate, with respect to which variable should we differentiate this function?

OpenStudy (anonymous):

\[(2nb^{rx}+n)^k\] is the equation that has to be differentiated

OpenStudy (amistre64):

you mean find a derivative?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

if so; assume they are all functions of the same variable; and sort it out when you know more info :)

OpenStudy (amistre64):

of course i got no idea how to recall to do an x^x type deal; but if you can remember it :) otherwise ill assume k to be constant \[D(2nb^{rx}+n)^k=k(2nb^{rx}+n)^{(k-1)}*D(2nb^{rx}+n)\] \[=k(2nb^{rx}+n)^{(k-1)}*(D(2nb^{rx})+D(n))\] \[=k(2nb^{rx}+n)^{(k-1)}*(D(2nb^{rx})+n')\] \[=k(2nb^{rx}+n)^{(k-1)}*(D(2n)b^{rx}+2nD(b^{rx})+n')\] \[=k(2nb^{rx}+n)^{(k-1)}*(2n'b^{rx}+2nD(b^{rx})+n')\] something similar to that is what id assume

OpenStudy (anonymous):

Basically you have taken the derivative with respect to k

OpenStudy (amistre64):

y = x^x Ly = Lx^x Ly = x Lx y'/y = x' Lx+x L'x y'/y = Lx+x/x y'/y = Lx+1 y' = y(Lx+1) y' = x^x (Lx+1) y' = x^x Lx + x^x might be what I was forgetting

OpenStudy (amistre64):

well, I kept k constant and took the derivative of what I could with respect to say: time

OpenStudy (amistre64):

\[(2(n(t))(b(t))^{(r(t)x(t))}+n(t))^k\] somehting like this

OpenStudy (amistre64):

but thats only becasue I have no idea what the question is in context really

OpenStudy (anonymous):

The question just says differentiate. Thats it. Doesn't even ask with respect to which variable should the derivative be taken. Because the derivative changes with each variable

OpenStudy (anonymous):

For example, if the derivative is taken with respect to x, the derivative of the function has a different answer... if it is taken with respect to n, the answer changes accordingly.

OpenStudy (amistre64):

if we are to take partial derivaitves; meaning that all the variables are independant ... then yes

OpenStudy (amistre64):

i was thinking more of an implicit version